Two cars are 261 miles apart, they travel towards each other and meet in 3 Hrs. 1 car travels 1 MPH faster than the other. What is average speed of each car?
I believe I have this set up correctly but instructor marked it incorrect.
D/t=r
261/3=r
87=r of one car
R+1=88mph of second car
Is there anything I am leaving our?

Answer by stanbon(75887)

Two cars are 261 miles apart, they travel towards each other and meet in 3 Hrs. 1 car travels 1 MPH faster than the other. What is average speed of each car?
----
Car 1 DATA:
time = 3 hrs ; distance = x miles ; rate = x/3 mph
---
Car 2 DATA:
time = 3 hrs ; distance = 261-x miles ; rate = (261-x)/3 mph
----
Equation:
Car 1 rate - Car 2 rate = 1 mph
x/3 - (261-x)/3 = 1
----
Multiply thru by 3 to get:
x -(261-x) = 3
2x = 3+261
2x = 264
x = 132
261-x = 129
--
Car 1 rate: 132/3 = 44 mph
Car 2 rate: 129/3 = 43 mph
------------------------------
Cheers,
Stan H.
===================
Car 2 rate: